OpenStudy (anonymous):

will fan & medal :) (dont get what they are asking) Which of the following is not equivalent to the formula for the circumference of a circle, C = 2pir ? C over r equals 2 times pi C over pi equals 2 times r C over the quantity of r times pi equals 2 Cr = 2pi

OpenStudy (anonymous):

@decentnabeel

OpenStudy (decentnabeel):

A)C over r equals 2 times pi C/r = 2*pi because we divided both sides by r B)C over pi equals 2 times r C/pi = 2* r because we divided both sides by pi C)C over the quantity of r times pi equals 2 C/(pi*r) = 2 because we divded both sides by pi * r D)Cr = 2pi this is WRONG BECAUSE C = 2*pi * r you cant do any math to get C*r without an r on the other side too

OpenStudy (anonymous):

OHHHHHHH that was easy ... ughh idk why i didnt try harder ... thank you so much

OpenStudy (decentnabeel):

wellcome